OpenStudy (anonymous):

SUBTRACT THE POLYNOMIALS (7xy^3+5x^3y^4)-(4xy^3-2x^3y^4)

OpenStudy (anonymous):

DO NOT FACTOR

OpenStudy (anonymous):

simplify answer

geerky42 (geerky42):

Just subtract the like terms.

OpenStudy (anonymous):

walk me through it please..!!!

OpenStudy (anonymous):

..?

geerky42 (geerky42):

(7xy³ + 5x³y⁴) - (4xy³ - 2x³y⁴) (7xy³ + 5x³y⁴)+(-1)(4xy³ - 2x³y⁴) 7xy³ + 5x³y⁴ - 4xy³ + 2x³y⁴ Now combine like terms.

OpenStudy (anonymous):

ugh i cant . i ve been stuck for about an hour

OpenStudy (anonymous):

3xy^3?

geerky42 (geerky42):

Yes, because 7 - 4 = 3.

OpenStudy (anonymous):

3x^3

OpenStudy (anonymous):

and the last one?

geerky42 (geerky42):

7xy³ - 4xy³ (7 - 4)(xy³) 3xy³

geerky42 (geerky42):

For last one, just do what I showed you.

geerky42 (geerky42):

5x³y⁴ + 2x³y⁴ ???

geerky42 (geerky42):

Factor x³y⁴ out.

OpenStudy (anonymous):

7x^3y^4?

geerky42 (geerky42):

Yes!

OpenStudy (anonymous):

great picture i love team fortress lol

OpenStudy (anonymous):

so all together 3xy^3+7x^3y^4

geerky42 (geerky42):

Yes.

geerky42 (geerky42):

Is this clear?

OpenStudy (anonymous):

now it is thanks so much

geerky42 (geerky42):

Glad I helped.

OpenStudy (anonymous):

can you help with one more please

OpenStudy (anonymous):

??

geerky42 (geerky42):

Sure.

OpenStudy (anonymous):

i need to combine like terms with 3a*9a^2+3a*6a+3a*3+3*9a^2+3*6a+3*3

OpenStudy (anonymous):

i know the first one is 27a^3

geerky42 (geerky42):

What about second one? (3a*6a)

OpenStudy (anonymous):

18a?

OpenStudy (anonymous):

^2?

geerky42 (geerky42):

a·a = a² so 3a·6a = 3·6·a·a = 18a² Yes. 18a²

geerky42 (geerky42):

What don't you understand? I think you can handle it.

OpenStudy (anonymous):

thats where i get stuck. 27a^3+18A^2+ ?? 9A? +3????? AHHHH

geerky42 (geerky42):

Next one. 3a*3 What is it?

OpenStudy (anonymous):

9a?

geerky42 (geerky42):

Yes. What about 3*9a²?

OpenStudy (anonymous):

27a^2

OpenStudy (anonymous):

?

geerky42 (geerky42):

Yes. 3*6a = ?

OpenStudy (anonymous):

18a?

geerky42 (geerky42):

Yes! Last one. 3*3 = ?

OpenStudy (anonymous):

9!

geerky42 (geerky42):

Excellent! Now you go from 3a*9a² + 3a*6a + 3a*3 + 3*9a² + 3*6a + 3*3 To 27a³ + 18a² + 9a + 27a² + 18a + 9. Now it is cleaner. Now all you have to do is combining like terms. Do you know what to do?

OpenStudy (anonymous):

...not really lets keep it going i get it so far

geerky42 (geerky42):

Which term has a²? What about a?

OpenStudy (anonymous):

is this right? 27a^3+45a^2+27a+9 ???

geerky42 (geerky42):

18a² and 27a² both share a². These terms are called like terms. 9a and 18a are like terms. You can combine them together by add 'coefficient' (18 and 27). Is this clear? Yes it's 27a³ + 45a² + 27a + 9.

OpenStudy (anonymous):

So i did it RIGHTTT!?

geerky42 (geerky42):

Yeah.

geerky42 (geerky42):

Does this make sense now?

OpenStudy (anonymous):

it told me i got it wrong.. =(

OpenStudy (anonymous):

apparently the correct answer was 27a^3+81a^2+81a+27

geerky42 (geerky42):

Are you sure you didn't typo? 3a*9a² + 3a*6a + 3a*3 + 3*9a² + 3*6a + 3*3